A man is six times as old as his son. In 7 1/2 years, he will be just three times as old as his son will be then. How old is the man?
Answer by ankor@dixie-net.com(22740)

m = man's age
s = son's age
:
Write an equation for each statement
:
A man is six times as old as his son.
m = 6s
:
In 7 1/2 years, he will be just three times as old as his son will be then.
m + 7.5 = 3(s+7.5)
m + 7.5 = 3s + 22.5
m = 3s + 22.5 - 7.5
m = 3s + 15
replace m with 6s
6s - 3s = 15
3s = 15
s = 15/3
s = 5 yrs is son's age
then
6(5) = 30 yrs is the man's age
:
:
Check solutions in the 2nd statement
30 + 7.5 = 3(5 + 7.5)
37.5 = 3(12.5)